1. Write the point-slope form of the equation passing through (5, -1) with a slope of 6.
Answer by friesr(113)

y = mx + b is the formula
-
the give you slope "m" of 6
-
Solve for b
-1 = 6(5) + b
---
rewrite
y = 6x + b replacing b with the anser you get above.